# Regularized least squares learning with heavy-tailed noise is minimax optimal

### v2

@@ 4 Capacity dependent bound and optimal rates
We now improve the results from the previous section and give an excess risk bound that involves the effective dimension, which has been established as a central tool to quantify the algorithm-dependent capacity of the hypothesis space $\mathcal{H}$ relative to the distribution $\pi$ and regularization parameter $\alpha$ in order to derive risk bounds in regularized kernel-based learning under the assumptions of an eigenvalue decay of $\mathcal{C}_{\pi}$ [31, 3, 9, 10].
@@
We now assume the standard polynomial eigenvalue decay of $\mathcal{C}_{\pi}$ [3, 9, 11].
@@
The constant $c$ is made explicit in the proof. The key idea builds upon the original work of [3]. In particular, we incorporate the effective dimension into the Fuk–Nagaev inequality as a proxy for the $q$-th absolute moment appearing in the term $\eta(\delta,n,\alpha)$, thereby generalizing the idea to use the effective dimension as a variance proxy in the Bernstein inequality.
@@
The rates provided by Corollary 4.4 match the rates from the literature derived for well-specified kernel ridge regression under the assumption of subexponential noise [3, 34]. In particular, these rates are known to be minimax optimal over the class of distributions satisfying (SRC), (EVD) and the Bernstein condition (1.3). Corollary 4.4 now proves that one can significantly relax the assumption of subexponential noise (1.3), as rate optimality is already achieved under the condition (MOM). Furthermore, the regularization schedule $\alpha(n,\delta)$ is the same as in the light-tailed setting—in particular, it does not depend on $q$.
@@ 5 Fuk–Nagaev inequality in Hilbert spaces
We discuss the central ingredient for the derivation of the previous results in more detail for convenience. We refer the reader to [24, 25] for the original work in the setting of real-valued random variables and [59] for a discussion of the involved constants. We present a sharpened version of a result due to [27, Theorem 3.5.1], which is formulated more generally in normed spaces, but exhibits an excess term that can be removed in the Hilbert space case. We provide the proof in Appendix C.
@@
for some constants $\sigma^{2}>0$, $Q>0$ and $q\geq 3$. Write $S_{n}:=\sum_{i=1}\xi_{i}$. Then there exist two constants $c_{1}>0$ and $c_{2}>0$ depending only on $q$ such that for every $t>0$, we have
@@
Both the subgaussian term and the polynomial term in the right hand side of the bound given by Proposition 5.1 can generally not be improved without additional assumptions. We repeat a similar argument as the one given in [60, Proposition 9], which is given in the context of linear processes. Let $\xi,\xi_{1},\dots,\xi_{n}$ be independent real-valued random variables drawn from a centered $t$-distribution with $q$ degrees of freedom, i.e. $\mathbb{E}[\xi^{q-c}]<\infty$ for all $0<c\leq q$ and let $\sigma^{2}:=\mathbb{E}[\xi^{2}]=q/(q-2)$. Then [25, Theorem 1.9] shows that we have
@@ Acknowledgements
Dimitri Meunier and Arthur Gretton are supported by the Gatsby Charitable Foundation. This work was partially funded by the German Federal Ministry of Research, Technology and Space as part of the 6G-CAMPUS project (grant no. 16KISK194).
@@
The appendices are organized as follows. In Appendix A, we report all proofs of the results in the main text of this paper. In Appendix B, we provide an additional excess risk bound for the polynomial confidence regime based on Proposition 3.1. We collect some tail bounds and concentration results in Appendix C. In particular, Section C.3 contains a proof of Proposition 5.1. In Appendix D, we recall additional miscellaneous inequalities required for our proofs.
Appendix A Proofs
@@
A.1.1 Bounding the variance $I_{\pi}\widehat{f}_{\alpha}-I_{\pi}f_{\alpha}$.
@@
$\mathbb{E}[\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})]=I_{\pi}^{*}(f_{\star}-I_{\pi}f_{\alpha})=\alpha f_{\alpha}.$ (A.1)
@@
$=(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}\mathopen{}\mathclose{{}\left\{\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})-I_{\pi}^{*}(f_{\star}-I_{\pi}f_{\alpha})+\frac{1}{n}\sum_{i=1}^{n}\phi(X_{i})\varepsilon_{i}}\right\}$ by (A.1)
@@
$\Delta_{1}=\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})-\mathbb{E}[\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})],\qquad\Delta_{2}=\frac{1}{n}\sum_{i=1}^{n}\phi(X_{i})\varepsilon_{i}\;.$ (A.2)
@@
$\leq 2\sqrt{\frac{2}{\alpha}}(\lVert\Delta_{1}\rVert_{\mathcal{H}}+\lVert\Delta_{2}\rVert_{\mathcal{H}}),$ (A.3)
with probability at least $1-\delta$, where ${\mathcal{B}}(n,\delta,\alpha)$ is defined in (C.4) and where we apply Lemma C.10 in the first inequality, Corollary C.7 in the second inequality, and additionally the fact that we have $\lVert(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1/2}\rVert_{\mathcal{L}(\mathcal{H})}\leq\alpha^{-1/2}$ in the last inequality.
@@
We introduce the independent and identically distributed random variables $\xi_{i}:=(f_{\star}(X_{i})-f_{\alpha}(X_{i}))\phi(X_{i})\in\mathcal{H}$ and note that $\Delta_{1}:=\frac{1}{n}\sum_{i=1}^{n}(\xi_{i}-\mathbb{E}[\xi_{i}])$ and proceed similarly as in the proof of [4, Theorem 1 & Theorem 5].
@@
Furthermore, by [61, Proposition 3.3] and [11, Lemma 25] we see
@@
Applying Proposition C.2 to $\Delta_{1}=\frac{1}{n}\sum_{i=1}^{n}(\xi_{i}-\mathbb{E}[\xi_{i}])$, with probability at least $1-\delta$ for for any $01>\delta>$ 0, we have
@@
$\leq\tilde{C}_{\diamond}\cdot\mathopen{}\mathclose{{}\left(\frac{\log(2/\delta)}{n}+\sqrt{\frac{\alpha^{2\min\{\nu,1\}}\log(2/\delta)}{n}}}\right),$ (A.4)
@@
$\tilde{C}_{\diamond}=2\kappa\cdot\max\{\lVert f_{\star}\rVert_{L^{\infty}(\pi)}+R\cdot\kappa^{2\nu},R\}.$ (A.5)
@@
$\mathopen{}\mathclose{{}\left\|\Delta_{2}}\right\|_{\mathcal{H}}\leq\max\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(\frac{2c_{1}\kappa^{q}Q}{\delta n^{q-1}}}\right)^{1/q},\kappa\sigma\;\sqrt{\frac{\log(2c_{1}/\delta)}{c_{2}n}}}\right\}$ (A.6)
@@
By the union bound, the three individual confidence bounds (A.3), (A.1.1) and (A.6) and hold simultaneously with probability at least $1-\delta$. We have
@@
$C_{\diamond}=2\sqrt{2}\max\{\tilde{C}_{\diamond},\kappa\cdot\max\{(6c_{1})^{1/q},1/\sqrt{c_{2}}\}\}.$ (A.7)
A.1.2 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
We apply the standard theory of source conditions in inverse problems as discussed for example by [55]. In particular, under the $L^{2}(\pi)$-source condition given by (SRC), we have
@@ A.1.2 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
see e.g. [61, Proposition 3.3]. Combining the variance bound and the bias bound yields the result.
A.2 Proof of Corollary 3.2
@@
$\alpha^{\min\{\nu,1\}}\sqrt{n}\geq\sqrt{\log(6/\delta)}$ (A.8)
@@
$\alpha\geq c_{4}\;\mathopen{}\mathclose{{}\left(\frac{\log(6c_{1}/\delta)}{n}}\right)^{\frac{1}{2\min\{\nu,1\}+1}},$ (A.9)
@@
with confidence $1-\delta$ and $\tilde{c}_{1}=2c_{4}^{\min\{\nu,1\}}$. In order for this bound to hold, the regularization strength has to fulfill condition (3.1), (A.8) and (A.9), that is,
$\alpha\geq\max\mathopen{}\mathclose{{}\left\{c_{4}\mathopen{}\mathclose{{}\left(\frac{\log(6c_{1}/\delta)}{n}}\right)^{\frac{1}{2\min\{\nu,1\}+1}},\mathopen{}\mathclose{{}\left(\frac{\log(6/\delta)}{n}}\right)^{\frac{1}{2\min\{\nu,1\}}},C_{\kappa}\frac{\log(6/\delta)}{\sqrt{n}}}\right\}$ (A.10)
@@
satisfies (A.10).
A.3 Proof of Proposition 4.3
@@
A.3.1 Bounding the variance $I_{\pi}\widehat{f}_{\alpha}-I_{\pi}f_{\alpha}$
Following [62, Appendix C], we further decompose the variance term in $L^{2}(\pi)$ as

### v3

@@ 4 Capacity dependent bound and optimal rates
We now improve the results from the previous section and give an excess risk bound that involves the effective dimension, which has been established as a central tool to quantify the algorithm-dependent capacity of the hypothesis space $\mathcal{H}$ relative to the distribution $\pi$ and regularization parameter $\alpha$ in order to derive risk bounds in regularized kernel-based learning under the assumptions of an eigenvalue decay of $\mathcal{C}_{\pi}$ [zhang05learning, caponnetto2007optimal, Muecke18, Lin2020].
@@
We now assume the standard polynomial eigenvalue decay of $\mathcal{C}_{\pi}$ [caponnetto2007optimal, Muecke18, fischer2020sobolev].
@@
The constant $c$ is made explicit in the proof. The key idea builds upon the original work of [caponnetto2007optimal]. In particular, we incorporate the effective dimension into the Fuk–Nagaev inequality as a proxy for the $q$-th absolute moment appearing in the term $\eta(\delta,n,\alpha)$, thereby generalizing the idea to use the effective dimension as a variance proxy in the Bernstein inequality.
@@
The rates provided by Corollary 4.4 match the rates from the literature derived for well-specified kernel ridge regression under the assumption of subexponential noise [caponnetto2007optimal, Steinwart09Optimal]. In particular, these rates are known to be minimax optimal over the class of distributions satisfying (SRC), (EVD) and the Bernstein condition (1.3). Corollary 4.4 now proves that one can significantly relax the assumption of subexponential noise (1.3), as rate optimality is already achieved under the condition (MOM). Furthermore, the regularization schedule $\alpha(n,\delta)$ is the same as in the light-tailed setting—in particular, it does not depend on $q$.
@@ 5 Fuk–Nagaev inequality in Hilbert spaces
We discuss the central ingredient for the derivation of the previous results in more detail for convenience. We refer the reader to [Fuk1973, Nagaev1979] for the original work in the setting of real-valued random variables and [Rio2017] for a discussion of the involved constants. We present a sharpened version of a result due to [Yurinsky1995, Theorem 3.5.1], which is formulated more generally in normed spaces, but exhibits an excess term that can be removed in the Hilbert space case. We provide the proof in Appendix D. We also note that the proof of the result in [Yurinsky1995] is incomplete due to an inconsistent exponential moment bound. We address this issue by deriving an alternative bound.
@@
for some constants $\sigma^{2}>0$, $Q>0$ and $q\in\mathbb{N}$, $q\geq 3$. Write $S_{n}:=\sum_{i=1}\xi_{i}$. Then there exist two constants $c_{1}>0$ and $c_{2}>0$ depending only on $q$ such that for every $t>0$, we have
@@
Both the subgaussian term and the polynomial term in the right hand side of the bound given by Proposition 5.1 can generally not be improved without additional assumptions. We repeat a similar argument as the one given in [LouEtAl2022, Proposition 9], which is given in the context of linear processes. Let $\xi,\xi_{1},\dots,\xi_{n}$ be independent real-valued random variables drawn from a centered $t$-distribution with $q$ degrees of freedom, i.e. $\mathbb{E}[\xi^{q-c}]<\infty$ for all $0<c\leq q$ and let $\sigma^{2}:=\mathbb{E}[\xi^{2}]=q/(q-2)$. Then [Nagaev1979, Theorem 1.9] shows that we have
@@ Acknowledgements
Dimitri Meunier and Arthur Gretton are supported by the Gatsby Charitable Foundation. This work was partially funded by the German Federal Ministry of Research, Technology and Space as part of the 6G-CAMPUS project (grant no. 16KISK194). The authors express their gratitude to Christian Fiedler for pointing out inconsistensies in the proof of [Yurinsky1995, Lemma 3.5.1] and for his contribution to the corresponding corrections in Section D.3.1 and Section D.3.3.
@@
The appendices are organized as follows.
Appendix A contains a numerical experiment which confirms the behavior of the excess risk as described by the bounds provided in the main text in a very basic setting. In Appendix B, we report proofs of the results in the main text of this paper. In Appendix C, we provide an additional excess risk bound for the polynomial confidence regime based on Proposition 3.1. We collect some tail bounds and concentration results in Appendix D. In particular, Section D.3 contains a proof of Proposition 5.1. In Appendix E, we recall additional miscellaneous inequalities required for our proofs.
Appendix A Numerical experiment
We provide a basic numerical experiment [^3] We provide the source code on GitHub: https://github.com/mollenhauerm/krr-heavy-tailed. The experiment can be run on the CPU of any consumer laptop. showing that the confidence regimes are not only occurring in the upper bound on the excess risk given by Proposition 3.1, but are effectively reflected in the distribution of the excess risk in practice, even for very simple models.
We consider the input space $\mathcal{X}:=\mathbb{R}$ equipped with the RKHS $\mathcal{H}$ induced by the radial basis kernel $k(x_{1},x_{2}):=\exp(\tfrac{-|x_{1}-x_{2}|^{2}}{2})$ and define the target function $f_{\star}(x):=\sum_{i=1}^{5}a_{i}k(x_{i},x)\in\mathcal{H}$, $x\in\mathcal{X}$ for vectors $\mathbf{a}:=(2,-1,-3,1,2)$ and $\mathbf{x}:=(-4,-2-0,3,7)$. We define the covariate distribution $X\sim\pi:=\mathcal{N}(0,1)$ on $\mathcal{X}$ and generate independent observation pairs with a light-tailed noise distribution and a heavy-tailed distribution with identical variance based on
  1. (i)
    the light-tailed noise model given by
    $Y_{(\mathcal{N})}=f_{\star}(X)+\varepsilon_{(\mathcal{N})},$ (A.1)
    where the noise $\varepsilon_{(\mathcal{N})}\sim\mathcal{N}(0,\sigma^{2})$ follows a centered Gaussian distribution with variance $\sigma^{2}=3$.
  2. (ii)
    the heavy-tailed noise model with a finite number of higher moments given by
    $Y_{(t)}=f_{\star}(X)+\varepsilon_{(t)},$ (A.2)
    where the noise $\varepsilon_{(t)}\sim\mathbf{t}(0,\nu)$ follows a centered $t$-distribution with $\nu=3$ degrees of freedom and $\mathbb{E}[\varepsilon_{(t)}^{2}]=3$.
(a) Light-tailed noise
(b) Heavy-tailed noise
Figure A.1: Empirical approximations of the quantile function of the excess risk $\lVert I_{\pi}\widehat{f}_{\alpha}-f_{\star}\rVert_{L^{2}(\pi)}$ for (a) the light-tailed noise model given in (A.1) and (b) in the $t$-distributed noise model given by (A.2) for different choices of the regularization parameter $\alpha$ and fixed sample size $n=20$.
For both models, we compute $\widehat{f}_{\alpha}$ based on the generated sample pairs for a sample size $n=20$ and record the error $\lVert I_{\pi}\widehat{f}_{\alpha}-f_{\star}\rVert_{L^{2}(\pi)}$, which we approximate through Monte Carlo simulation by drawing samples from $\pi$. We perform the above computation for a selection of different regularization parameters $\alpha$, repeating each experiment across $10000$ random seeds (per model and choice of $\alpha$). Based on the recorded errors, we empirically approximate the quantile function of the distributions of $\lVert I_{\pi}\widehat{f}_{\alpha}-f_{\star}\rVert_{L^{2}(\pi)}$ given by
$Q(1-\delta):=\inf\mathopen{}\mathclose{{\left\{t\colon\mathbb{P}\mathopen{}\mathclose{{\left[\lVert I_{\pi}\widehat{f}_{\alpha}-f_{\star}\rVert_{L^{2}(\pi)}\leq t}}\right]\geq 1-\delta}}\right\},\quad\delta\in[0,1]$
The resulting approximated quantile functions are visualized in Figure A.1 for all choices of $\alpha$ and both models.
While the excess risk quantiles exhibit the same qualitative behaviour and are of the same order in both models for moderate confidence levels, the excess risk quantiles of the heavy-tailed model increase rapidly for a small regularization parameter beyond a high confidence level of $1-\delta\approx 0.95$. Consequently, the heavy-tailed model requires a stronger regularization than the light-tailed model in high confidence settings, as explained by the transition of the excess risk from a subgaussian confidence regime to a polynomial confidence regime in Section 3.
Appendix B Proofs
@@
B.1.1 Bounding the variance $I_{\pi}\widehat{f}_{\alpha}-I_{\pi}f_{\alpha}$.
@@
$\mathbb{E}[\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})]=I_{\pi}^{*}(f_{\star}-I_{\pi}f_{\alpha})=\alpha f_{\alpha}.$ (B.1)
@@
$=(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}\mathopen{}\mathclose{{\left\{\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})-I_{\pi}^{*}(f_{\star}-I_{\pi}f_{\alpha})+\frac{1}{n}\sum_{i=1}^{n}\phi(X_{i})\varepsilon_{i}}}\right\}$ by (B.1)
@@
$\Delta_{1}=\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})-\mathbb{E}[\widehat{I}_{\pi}^{*}(f_{\star}-\widehat{I}_{\pi}f_{\alpha})],\qquad\Delta_{2}=\frac{1}{n}\sum_{i=1}^{n}\phi(X_{i})\varepsilon_{i}\;.$ (B.2)
@@
$\leq 2\sqrt{\frac{2}{\alpha}}(\lVert\Delta_{1}\rVert_{\mathcal{H}}+\lVert\Delta_{2}\rVert_{\mathcal{H}}),$ (B.3)
with probability at least $1-\delta$, where ${\mathcal{B}}(n,\delta,\alpha)$ is defined in (D.8) and where we apply Lemma D.10 in the first inequality, Corollary D.7 in the second inequality, and additionally the fact that we have $\lVert(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1/2}\rVert_{\mathcal{L}(\mathcal{H})}\leq\alpha^{-1/2}$ in the last inequality.
@@
We introduce the independent and identically distributed random variables $\xi_{i}:=(f_{\star}(X_{i})-f_{\alpha}(X_{i}))\phi(X_{i})\in\mathcal{H}$ and note that $\Delta_{1}:=\frac{1}{n}\sum_{i=1}^{n}(\xi_{i}-\mathbb{E}[\xi_{i}])$ and proceed similarly as in the proof of [Smale2007, Theorem 1 & Theorem 5].
@@
Furthermore, by [DeVitoEtAl06, Proposition 3.3] and [fischer2020sobolev, Lemma 25] we see
@@
Applying Proposition D.2 to $\Delta_{1}=\frac{1}{n}\sum_{i=1}^{n}(\xi_{i}-\mathbb{E}[\xi_{i}])$, with probability at least $1-\delta$ for for any $1>\delta>0$, we have
@@
$\leq\tilde{C}_{\diamond}\cdot\mathopen{}\mathclose{{\left(\frac{\log(2/\delta)}{n}+\sqrt{\frac{\alpha^{2\min\{\nu,1\}}\log(2/\delta)}{n}}}}\right),$ (B.4)
@@
$\tilde{C}_{\diamond}=2\kappa\cdot\max\{\lVert f_{\star}\rVert_{L^{\infty}(\pi)}+R\cdot\kappa^{2\nu},R\}.$ (B.5)
@@
$\mathopen{}\mathclose{{\left\|\Delta_{2}}}\right\|_{\mathcal{H}}\leq\max\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(\frac{2c_{1}\kappa^{q}Q}{\delta n^{q-1}}}}\right)^{1/q},\kappa\sigma\;\sqrt{\frac{\log(2c_{1}/\delta)}{c_{2}n}}}}\right\}$ (B.6)
@@
By the union bound, the three individual confidence bounds (B.3), (B.1.1) and (B.6) and hold simultaneously with probability at least $1-\delta$. We have
@@
$C_{\diamond}=2\sqrt{2}\max\{\tilde{C}_{\diamond},\kappa\cdot\max\{(6c_{1})^{1/q},1/\sqrt{c_{2}}\}\}.$ (B.7)
B.1.2 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
We apply the standard theory of source conditions in inverse problems as discussed for example by [EHN96]. In particular, under the $L^{2}(\pi)$-source condition given by (SRC), we have
@@ B.1.2 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
see e.g. [DeVitoEtAl06, Proposition 3.3]. Combining the variance bound and the bias bound yields the result.
B.2 Proof of Corollary 3.2
@@
$\alpha^{\min\{\nu,1\}}\sqrt{n}\geq\sqrt{\log(6/\delta)}$ (B.8)
@@
$\alpha\geq c_{4}\;\mathopen{}\mathclose{{\left(\frac{\log(6c_{1}/\delta)}{n}}}\right)^{\frac{1}{2\min\{\nu,1\}+1}},$ (B.9)
@@
with confidence $1-\delta$ and $\tilde{c}_{1}=2c_{4}^{\min\{\nu,1\}}$. In order for this bound to hold, the regularization strength has to fulfill condition (3.1), (B.8) and (B.9), that is,
$\alpha\geq\max\mathopen{}\mathclose{{\left\{c_{4}\mathopen{}\mathclose{{\left(\frac{\log(6c_{1}/\delta)}{n}}}\right)^{\frac{1}{2\min\{\nu,1\}+1}},\mathopen{}\mathclose{{\left(\frac{\log(6/\delta)}{n}}}\right)^{\frac{1}{2\min\{\nu,1\}}},C_{\kappa}\frac{\log(6/\delta)}{\sqrt{n}}}}\right\}$ (B.10)
@@
satisfies (B.10).
B.3 Proof of Proposition 4.3
@@
B.3.1 Bounding the variance $I_{\pi}\widehat{f}_{\alpha}-I_{\pi}f_{\alpha}$
Following [mucke2019beating, Appendix C], we further decompose the variance term in $L^{2}(\pi)$ as